OpenStudy (anonymous):

How do I solve 3n^2(n^2+4)=20-2n^2

OpenStudy (anonymous):

I have this so far.\[3n^{2}(n^{2}+4)=2-2n^{2}\]\[3n^{4}+12n^{2}=20-n^{2}\]\[3n^{4}+14n^{2}-20=0\]

OpenStudy (imstuck):

The first thing to do is to distribute the 3n^2 into the parenthesis. That will give you the ability to move the n^2 terms around and add or subtract them as needed. So let's do that: